2792000217:

If g(x) = 3(x − 2), find the value of x if g(x) = 6. (2 points) 12 2.6 8 4

mxddi3:

so g(x)= 3(x-2) and g(x)=6. so first step rewrite it. \[g(6)=3(6-2)\] does this make sense so far?

2792000217:

Yes

mxddi3:

ok, so, now you distribute. what is 3(6-2) ?

2792000217:

12

mxddi3:

yes, good job. sooo, g(6)=12

mxddi3:

does this make sense to you?

2792000217:

Kinda it takes me a little to latch on to it I hope u understand

mxddi3:

yes, i do. Is there anything I can do to help you further understand?

2792000217:

Can you help me go through the steps

mxddi3:

yes. ok so start from the beginning.

mxddi3:

it says, g(x)=3(x-2) when g(x)=6. what this is saying is that x=6. so since you are given the variable, you replace the x with the 6. does this make sense or do you need me to explain it different?

2792000217:

Perfect since thx So wait would my answer be 12

2792000217:

*sinse

mxddi3:

yes, it would :) do you understand how it's 12?

2792000217:

*sense sorry im brain dead today

2792000217:

yes i do thank you for ur help

mxddi3:

you're welcome, glad to help!

2792000217:

If i need ur help will u help me again

mxddi3:

ofc!!! just ping me if I don't see it withing the first few minutes that you post!

mxddi3:

within*

2792000217:

ok

mxddi3:

:) you can close the post now !!